# Jetfire: Efficient and Accurate Transformer Pretraining with INT8 Data Flow and Per-Block Quantization

### v1

@@ Abstract
Pretraining transformers are generally time-consuming. Fully quantized training (FQT) is a promising approach to speed up pretraining. However, most FQT methods adopt a quantize-compute-dequantize procedure, which often leads to suboptimal speedup and significant performance degradation when used in transformers due to the high memory access overheads and low-precision computations. In this work, we propose Jetfire, an efficient and accurate INT8 training method specific to transformers. Our method features an INT8 data flow to optimize memory access and a per-block quantization method to maintain the accuracy of pretrained transformers. Extensive experiments demonstrate that our INT8 FQT method achieves comparable accuracy to the FP16 training baseline and outperforms the existing INT8 training works for transformers. Moreover, for a standard transformer block, our method offers an end-to-end training speedup of 1.42x and a 1.49x memory reduction compared to the FP16 baseline.
@@ 2 Related Work
For INT8 training, the majority of the work focuses on quantization of CNNs (Zhu et al., 2020; Zhao et al., 2021; Zhou et al., 2021). SwitchBack (Wortsman et al., 2023) introduces per-token quantization and successfully applies INT8 training to CLIP models for the first time, but still leaves the calculation of weight gradient in FP. For LLM pre-training, per-token quantization still results in significant accuracy loss due to channel-wise outliers.
@@ 7.1 Settings
We quantize all of the linear layers in the MLP and attention module and non-linear layers (including GELU, LayerNorm and Dropout) to INT8, and leave multi-head attention in FP16 by employing FlashAttention (Dao et al., 2022). We compare our method with floating point training baseline (denoted as FP), per-tensor quantization, and SwitchBack (Wortsman et al., 2023)). We do not compare with FP8 training algorithms as they require specialized Hopper architecture GPU to run, making them less accessible. We emphasize that only our method adopts an INT8 data flow and quantizes non-linear layers.
@@ 7.2 Converged Model Accuracy
Image Classification
We do pretraining for Deit-Tiny, Deit-Small, and Deit-Base (Touvron et al., 2021) model on ImageNet1K (Deng et al., 2009) for 90 epochs based on facebook research’s recipe  [^2] https://github.com/facebookresearch/deit. Results are listed on Table 4. In all experiments, Our method has less than $0.1\%$ accuracy degradation compared with the floating-point baseline, and for Deit-base, our method shows $0.4\%$ improvement. For Deit-tiny and Deit-small models, Switchback has over $0.5\%$ accuracy degradation, and per-tensor quantization does not converge. This indicates that our method can accurately quantize channel-wise outliers.
@@ 7.4 Operator and End-to-End experiments
We test the speedup of our custom linear layer. We analyzed the time consumption of each component in forward and backward passes and compared the speed of our implementation with FP16 and SwitchBack linear layers. The results are shown in Fig. 8. Our MM operator provides about 60% speed improvement compared to FP16. Other overhead components like quantizing and reshaping have a relatively minor impact. Our method achieves 40% overall speedup (forward + backward), which is comparable to the acceleration result of SwitchBack, where SwitchBack leaves the calculation of weight gradient in FP. The speedup becomes larger when the matrix size increases since the overhead proportion decreases.
@@
We experimented with GPT2 models and varied the network hidden size to show the end-to-end speedup for our Jetfire method over PyTorch’s FP16 training. We integrated all linear and non-linear operators and reported the speedup of a transformer layer. We compared the forward, backward, and overall runtime speedup with the SwitchBack layer. Results in Table 5 showed that our method achieved comparable or improved acceleration compared to SwitchBack. This is primarily because our linear operators in backpropagation are faster than SwitchBack, and we can accelerate all of the non-linear operators in both forward and backward propagation.
Table 5: Acceleration ratios for End-to-end comparison (SB refers to SwitchBack basic version)
@@
Table 6: Activation memory reduction ratios for End-to-end comparison (SB refers to SwitchBack Memeory Efficient version)
@@ Acknowledgements
The authors would like to thank Bingrui Li, Ziteng Wang, Jiayi Zhong, Cheng Lu for the helpful discussions.
9 Broader Impact

### v2

@@ Abstract
Pretraining transformers are generally time-consuming. Fully quantized training (FQT) is a promising approach to speed up pretraining. However, most FQT methods adopt a quantize-compute-dequantize procedure, which often leads to suboptimal speedup and significant performance degradation when used in transformers due to the high memory access overheads and low-precision computations.
In this work, we propose Jetfire, an efficient and accurate INT8 training method specific to transformers. Our method features an INT8 data flow to optimize memory access and a per-block quantization method to maintain the accuracy of pretrained transformers. Extensive experiments demonstrate that our INT8 FQT method achieves comparable accuracy to the FP16 training baseline and outperforms the existing INT8 training works for transformers. Moreover, for a standard transformer block, our method offers an end-to-end training speedup of 1.42x and a 1.49x memory reduction compared to the FP16 baseline. Our code is open sourced at https://github.com/thu-ml/Jetfire-INT8Training.
@@ 2 Related Work
For INT8 training, the majority of the work focuses on quantization of CNNs (Zhu et al., 2020; Zhao et al., 2021; Zhou et al., 2021). SwitchBack (Wortsman et al., 2023) introduces per-token quantization and successfully applies INT8 training to CLIP models for the first time, but still leaves the calculation of weight gradient in FP. To be more specific, in the forward process of $\boldsymbol{\mathbf{Y}}=\boldsymbol{\mathbf{X}}\boldsymbol{\mathbf{W}}^{\top}$, they apply per-token quantization for $\boldsymbol{\mathbf{X}}$ and per-channel quantization for $\boldsymbol{\mathbf{W}}^{\top}$. In the backward process, for $\nabla\boldsymbol{\mathbf{X}}=\nabla\boldsymbol{\mathbf{Y}}\boldsymbol{\mathbf{W}}$, they apply per-token quantization for $\nabla\boldsymbol{\mathbf{Y}}$ and per-channel quantization for $\boldsymbol{\mathbf{W}}$, and leave the calculation of $\nabla\boldsymbol{\mathbf{W}}=\nabla\boldsymbol{\mathbf{Y}}^{\top}\boldsymbol{\mathbf{X}}$ in full precision. For LLM pre-training, per-token quantization still results in significant accuracy loss due to
@@ 7 Experiments
Model Swin-tiny Swin-small Swin-base
FP 77.55 80.39 80.45
Jetfire 77.51 80.39 80.37
ViT-base ViT-large
FP 83.45 85.72
Jetfire 83.48 85.67
Table 5: Comparison of FP and Jetfire
@@ 7.1 Settings
We quantize all of the linear layers in the MLP and attention module and non-linear layers (including GELU, LayerNorm and Dropout) to INT8, and leave multi-head attention in FP16 by employing FlashAttention (Dao et al., 2022). The master copy of the weights is kept in FP32. We quantize linear layers’ weights to INT8 prior to each matmul, but leave layernorm’s weight and bias to floating-point since they are relatively small. We compare our method with floating point training baseline (denoted as FP), per-tensor quantization, and SwitchBack (Wortsman et al., 2023)). We do not compare with FP8 training algorithms as they require specialized Hopper architecture GPU to run, making them less accessible. We emphasize that only our method adopts an INT8 data flow and quantizes non-linear layers.
@@ 7.2 Converged Model Accuracy
Image Classification - Deit
We do pretraining for Deit-Tiny, Deit-Small, and Deit-Base (Touvron et al., 2021) model on ImageNet1K (Deng et al., 2009) for 90 epochs based on facebook research’s recipe  [^2] https://github.com/facebookresearch/deit. Results are listed on Table 4. In all experiments, Our method has less than $0.1\%$ accuracy degradation compared with the floating-point baseline, and for Deit-base, our method shows $0.4\%$ improvement. For Deit-tiny and Deit-small models, Switchback has over $0.5\%$ accuracy degradation, and per-tensor quantization does not converge. This indicates that our method can accurately quantize channel-wise outliers. Comparison with more baselines (wang2023GDA; zhao2021DAQ) can be found in Appendix C
Image Classification - Swin Transformers and ViT
We do pretraining for Swin-Transformers(Swin-tiny, Swin-small, Swin-base) for 90 epochs and fine-tuned ViT(ViT-base, ViT-large) for 100 epochs without pre-training (MAE includes pretraining and finetuning) on ImageNet1K. We adopt the official training recipe  [^3] https://github.com/microsoft/Swin-Transformer [^4] https://github.com/facebookresearch/mae?tab=readme-ov-file and default hyperparameters, and only compare with the full precision training baseline. The results are shown in Figure 5. In all of the experiments, our method achieves less than 0.1% accuracy degradation, which proves the accuracy of our method.
@@ 7.4 Operator and End-to-End experiments
We test the speedup of our custom linear layer on RTX 4090. We analyzed the time consumption of each component in forward and backward passes and compared the speed of our implementation with FP16 and SwitchBack linear layers. The results are shown in Fig. 8. Our MM operator provides about 60% speed improvement compared to FP16. Other overhead components like quantizing and reshaping have a relatively minor impact. Our method achieves 40% overall speedup (forward + backward), which is comparable to the acceleration result of SwitchBack, where SwitchBack leaves the calculation of weight gradient in FP. The speedup becomes larger when the matrix size increases since the overhead proportion decreases, which is demonstrated in Table 10. Acceleration results on RTX 3090 can be found in Appendix D.2.
@@
We experimented with GPT2 models and varied the network hidden size to show the end-to-end speedup for our Jetfire method over PyTorch’s FP16 training on RTX 4090. We integrated all linear and non-linear operators and reported the speedup of a transformer layer. We compared the forward, backward, and overall runtime speedup with the SwitchBack layer. Results in Table 6 showed that our method achieved comparable or improved acceleration compared to SwitchBack. This is primarily because our linear operators in backpropagation are faster than SwitchBack, and we can accelerate all of the non-linear operators in both forward and backward propagation. Acceleration results on RTX 3090 can be found in Appendix 11.
Table 6: Acceleration ratios for End-to-end comparison (SB refers to SwitchBack basic version) on GPT2 model.
@@
Table 7: Activation memory reduction ratios for End-to-end comparison (SB refers to SwitchBack Memeory Efficient version) on GPT2 model.
@@ Acknowledgements
The authors would like to thank Bingrui Li, Ziteng Wang, Jiayi Zhong, Cheng Lu for the helpful discussions. This work was supported by the National Science and Technology Major Project (2021ZD0110502), NSFC Projects (Nos. 62376131, 62061136001, 62106123, 62076147, U19A2081, 61972224), Tsinghua Institute for Guo Qiang, and the High Performance Computing Center, Tsinghua University. J.Z is also supported by the XPlorer Prize.
Impact Statement
@@ Appendix B Detailed Results of GLUE Fine-Tuning Test
Appendix C Comparisons with methods targeting CNNs
In this section, We tested two INT8 training for CNN models (wang2023GDA; zhao2021DAQ) on the DeiT pretraining experiment. As reported in Table 9, both of them showed significant accuracy degradation. This indicates that these methods are not sufficient to be applied to transformer models.
Model FP SwitchBack Per-tensor Jetfire GDA DAQ
Deit-tiny 64.08 63.55 - 63.95 62.14 61.80
Deit-small 73.43 72.80 - 73.31 70.98 70.66
Deit-base 75.67 75.62 - 76.03 73.06 72.40
Table 9: Comparison of different methods on various Deit models
Appendix D Acceleration Experiments
D.1 Overhead portion in Linear Layer
We tested the percentage of time taken by all quantization, dequantization, transpose, and other overhead processes during the forward and backward passes in a linear layer. We find that in Table 10, the relative overhead from quantization/dequantization diminishes with increasing model size, leading to more significant speed improvements.
Sequence Length
Batch Size Place 1024 2048 4096 6144 8192
4 forward 44.5 22.1 10.2 6.8 5.5
4 backward 30.4 29.0 16.6 12.2 9.2
4 overall 34.5 26.7 14.5 10.3 7.9
16 forward 23.1 8.1 2.7 1.7 1.3
16 backward 24.6 26.4 8.7 7.3 6.2
16 overall 23.7 18.3 11.5 9.9 8.9
Table 10: Percentage of overhead in a linear layer.
D.2 Acceleration result on other hardware
Besides RTX 4090, we tested our linear operator and end-to-end speed up result on the RTX 3090 GPUs, as reported in Table 11. The results indicate that our method can achieves significant speedups on multiple kinds of GPUs.
Size Settings Fwd-SB Fwd-Ours Bwd-SB Bwd-Ours All-SB All-Ours
Linear Layer, C=D=2048 1.60 1.53 1.38 1.31 1.45 1.38
Linear Layer, C=D=4096 2.43 1.87 1.37 1.48 1.62 1.60
Linear Layer, C=D=8192 2.56 1.70 1.24 1.40 1.51 1.49
End-to-End, hidden=1024 1.08 0.94 1.08 1.10 1.08 1.05
End-to-End, hidden=2048 1.34 1.18 1.15 1.36 1.21 1.29
End-to-End, hidden=4096 1.27 1.23 1.18 1.37 1.24 1.32
Table 11: Speed up result on the RTX 3090 GPUs. SB refers to SwitchBack, Ours refers to Jetfire.